August 22, 2006
- At the Dental School, the pipe fabrication for the medical gas systems continues.
- Oxygen piping is complete, while the NO2 is approximately 50% complete.
- The project to remove the flood damaged medical records is progressing well.
- The hoist cables, governor, and the travel cables have been replaced on the freight elevator #5.
- The hoist and travel cables were replaced on elevators #1, 2, and 4.
- The contractor making the temporary repairs at 1542 Tulane received the 13.8 K Volt switch gear was delivered today and it was moved into the basement. The installation can now begin.
- The contractor repairing the windows on the downtown campus addressed damage to parts of the metal curtains walls on the Nursing/Allied Health Building and the CSRB.